Compute the population proportion interval estimate given n, p, and the confidence level


a. Confidence Level= 99%, p=0.4, n=40

b. Confidence Level= 90%, p=0.15, n=55


1
Expert's answer
2022-06-02T13:54:38-0400

a. The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} = 2.5758."

The corresponding confidence interval is computed as shown below:



"CI(Proportion)=(\hat{p}-z_c\times \sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},""\hat{p}+z_c\times \sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})""=(0.4-2.5758\times \sqrt{\dfrac{0.4(1-0.4)}{40}},""0.4+2.5758\times \sqrt{\dfrac{0.4(1-0.4)}{40}})""=(0.20048, 0.59952)"

Therefore, based on the data provided, the 99% confidence interval for the population proportion is "0.20048 < p < 0.59952," which indicates that we are 99% confident that the true population proportion "p" is contained by the interval "(0.20048, 0.59952)."


b. The critical value for "\alpha = 0.10" is "z_c = z_{1-\alpha/2} =1.6449."

The corresponding confidence interval is computed as shown below:



"CI(Proportion)=(\hat{p}-z_c\times \sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},""\hat{p}+z_c\times \sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})""=(0.15-1.6449\times \sqrt{\dfrac{0.15(1-0.15)}{55}},""0.15+1.6449\times \sqrt{\dfrac{0.15(1-0.15)}{55}})""=(0.06975, 0.23025)"

Therefore, based on the data provided, the 99% confidence interval for the population proportion is "0.06975 < p < 0.23025," which indicates that we are 90% confident that the true population proportion "p" is contained by the interval "(0.06975, 0.23025)."